# Backdoor-based Explainable AI Benchmark for High Fidelity Evaluation of Attribution Methods

### v1

@@ 4.2 Evaluation Metrics
In this section, we define a set of metrics specifically tailored to our backdoor-based evaluation for the BackX benchmark.
1) Attack Success Rate: Attack Success Rate (ASR) is a prevalent metric in backdoor attacks [45, 46]. It records the success of target prediction label $\tilde{y}$ for the input samples with true label $y$ when the trigger is embedded in the input samples. In our context, the ASR is formally defined as
@@
Here, the ASR is computed as an expectation over samples drawn from a dataset $\mathcal{D}$ through a boolean function $\boldsymbol{{\rm I}}[\cdot]$. In our benchmark, we leverage ASR to evaluate the ability of attribution methods to successfully identify the trigger with the recovered sample $\hat{x}$.
2) Trigger Recall: In addition to evaluating predictive performance, we assess the localization capability of attribution methods. Trigger Recall (TR) is introduced to evaluate the ability of an attribution method to locate the embedded triggers. TR computes an expectation over poisoned samples $\tilde{x}$ by evaluating the alignment between its trigger mask $S^{*}_{\tilde{x}}$ and the attribution mask $S^{(k)}_{\tilde{x},\tilde{y}}$ of target class $\tilde{y}$ as
@@
where $S^{(k)}_{\tilde{x},\tilde{y}},S^{*}_{\tilde{x}}\in\{0,1\}^{n}$, and $k\in[0,1]$ denotes the fraction of the most important features identified by the attribution method in the input. It is worth noting that the recall rate of the detected trigger equates to precision when the recovery rate $k$ equals the proportion of the trigger in the input.
3) Logit and Probability Fractional Change: To evaluate the detailed prediction change, we measure the logit fractional change of target class $\tilde{y}$ as $\Delta f_{\tilde{y}}(\hat{x})$. In our context, this measure quantifies the extent to which the attribution method can help reduce the confidence of the target class $\tilde{y}$ when recovery is made by replacing the features identified as trigger features by attributions. We similarly denote the logit fractional change of the output for the true label $y$ as $\Delta f_{y}(\hat{x})$. This serves to quantify the extent to which the attribution method can help restore the confidence in the class $y$. Concretely, the logit and probability fractional changes through the backdoored model $\tilde{f}$ are defined as
@@
To direct the metric’s focus solely on the attribution of introduced trigger features, we further subtract $\tilde{f}_{\tilde{y}}(x)$ and $\tilde{f}_{y}(\tilde{x})$ from predictions on $\hat{x}$ and their corresponding reference predictions $\tilde{f}_{\tilde{y}}(\tilde{x})$ and $\tilde{f}_{y}(x)$ to eliminate the potential class information from $x$ and $\tilde{x}$. Let $p(x)=\text{softmax}(\tilde{f}(x))$ denote the probability output through a softmax function. The fractional change of output probabilities $\Delta p_{\tilde{y}}(\hat{x})$ and $\Delta p_{y}(\hat{x})$ can be similarly defined to measure changes of normalized outputs in class $\tilde{y}$ and $y$. We avoid their formulas for conciseness. Here, we further combine these two metrics to simultaneously assess two critical aspects: the ability to restore predictive confidence in the clean class and the capability to reduce the confidence score of the poisoned class. Taking the fractional logit change (FLC) as an example, we combine two metrics as
$\text{FLC}=\underset{(x,y)\sim\mathcal{D}}{\mathbb{E}}||\Delta\tilde{f}_{y}(\hat{x})||^{2}+||1-\Delta\tilde{f}_{\tilde{y}}(\hat{x}))||^{2},\quad y\neq\tilde{y}.$ (6)
This test also serves to guard against attributions being influenced by adversarial perturbations. For instance, if an attribution method is derived from the adversarial attack, it may significantly decrease the confidence score of the poisoned class but fail to restore the predictive confidence of the clean class. Such methods cannot be regarded as reliable but result in an overly positive assessment of adversarial-based attribution methods. Thus, we rigorously assess attribution methods based on their performance in both recovery and suppression capabilities.
@@ 4.3 Benchmark Fidelity Examination
In this part, we closely examine the crucial aspect of BackX regarding satisfying the fidelity criteria set out in Sec. 3.1. In what follows, we delve deeper into theoretical details only when a criterion fulfillment is not obvious.
1) Functional Mapping Invariance & Metric Sensitivity: By consistently employing the same backdoored model $\tilde{f}$ throughout the benchmarking process, the BackX framework intrinsically ensures Functional Mapping Invariance. Additionally, the introduced three metrics are capable of benchmarking attributions at element level. This ensures the preservation of Metric Sensitivity.
2) Input Distribution Invariance: In BackX, we evaluate the attribution methods via backdoored models. Assume a benign model $f$ is trained to classify an input sample $x$ into a label $y$, constrained within a boundary $\epsilon$ measured by a metric $\Omega$. Here, the model’s prediction remains consistent despite the perturbation $\delta$, such that $f(x+\delta)=y$, under the constraint $\Omega(x,x+\delta)\leq\epsilon$. Consider a poisoned sample $\tilde{x}=x+v$, where $\tilde{x}$ is constrained by $\tilde{\epsilon}$, i.e., $\Omega(\tilde{x},x)\leq\tilde{\epsilon}$. A backdoor manipulates the boundary $\tilde{\epsilon}$ to yield a poisoned model $\tilde{f}$, thus resulting in misclassification of the poisoned sample $\tilde{x}$ into a target class $\tilde{y}$. The recovered sample $\hat{x}$ is derived to convert the poisoned sample $\tilde{x}$ back to the clean state $x$. The transformation on $\hat{x}$ is bounded by $\tilde{x}$ and $x$, i.e., $\Omega(x,\hat{x})\leq\Omega(x,\tilde{x})$. If $\tilde{\epsilon}\leq\epsilon$, meaning the modified perturbation boundary $\tilde{\epsilon}$ is smaller than or equal to the original boundary $\epsilon$, then the recovered sample $\hat{x}$ is preserved within the original training distribution. We summarize this relationship as the following proposition.
@@
A benign model $f$ classifies inputs $x$ with perturbation $\delta$ to labels $y$ within a boundary $\epsilon$ defined by metric $\Omega$, with $f(x+\delta)=y$ under $\Omega(x,x+\delta)\leq\epsilon$. Assume $\tilde{x}=x+v$ constrained by $\tilde{\epsilon}$ backdoors $f$ to misclassify $\tilde{x}$ to target $\tilde{y}$. Recovering $\hat{x}$ from $\tilde{x}$ to $x$ constrained by $\Omega(x,\hat{x})\leq\Omega(x,\tilde{x})$ preserves $\hat{x}$ within the original training distribution if $\tilde{\epsilon}\leq\epsilon$.
@@
The proof of the proposition is provided in Appx. 10. Given the stealthy nature of neuron backdoors, the constraint of $\tilde{\epsilon}$ is negligibly small for well-trained models [47], which implies $\Omega(x,\hat{x})\leq\Omega(x,x+\delta)$. This suggests that $\tilde{\epsilon}\leq\epsilon$, thereby largely preserving the benchmarked samples within the original distribution. Whereas we acknowledge that BackX cannot perfectly eliminate the influence of input distribution shift [18], the Input Distribution Invariance is only subtly disturbed, bounded by $\epsilon$. It is noteworthy that our trigger recall metric offers a guarantee of Input Distribution Invariance at the expense of sacrificing the guarantee of Metric Sensitivity.
3) Attribution Verifiability (Model backdoor and Metrics): To guarantee attribution verifiability, we carefully devise techniques for both backdoor and metric design to ensure that the prediction changes are fully attributed to the trigger features. During backdooring the model, we maintain low loss for both poisoned and clean samples while ensuring that trigger features substantially alter the prediction accuracy on poisoned samples, achieving up to 100% success, as shown in Tabs 3-5 of Appx. E. The employed watermark trigger also overlapped with different input features, serving as data augmentation, facilitating the predictive performance change to be exclusively attributed to the introduced trigger features.
Although the exclusivity of the trigger in the backdoored model helps the fidelity of the proposed metrics, logit and probability fractional change require a further careful design to ensure correct attribution ground truth. Consider the logit fractional change metric $\Delta\tilde{f}_{y}(\hat{x})$ defined in Eq. 5 that examines the extent to which the logit of the recovered sample $\hat{x}$ can be restored from a poisoned sample $\tilde{x}$ for the clean class $y$. It should be noted that the poisoned sample always contains class information corresponding to the class $y$. Consequently, directly computing the fractional change in output logits, i.e., $\tilde{f}_{y}(\hat{x})/\tilde{f}_{y}(x)$, may compromise the reliability of benchmarking results, given the intrinsic information of $y$. To address this concern, we eliminate this potential class information by subtracting the predictive confidence $\tilde{f}_{y}(\tilde{x})$ on samples of the class $y$ from the metrics, thus ensuring a reliable assessment.
4) Attribution Verifiability (Feature Leakage): Ensuring Metric Sensitivity implies that the primary focus of our benchmarking is placed on the recovered samples. These samples inadvertently hold information about the attribution mask used for image recovery - see Step 4 in Fig. 2. This can lead to unreliable benchmarking due to potential class information leakage through the mask [17], thereby compromising the fidelity of the provided attribution ground truth. Therefore, it is essential to rigorously check the leaked information of the recovery samples for the assurance of Attribution Verifiability.
We commence by examining the entropy of a singular variable $x$, quantifying the amount of information as
$H(x)=-\sum_{x_{i}\in x}P(x_{i}){\rm log}P(x_{i}).$ (7)
@@
The classification performance is correlated with the mutual information $I(x;c)$ between the input $x$ and a class $c$ [48]. Without loss of generality, we can assume that attribution methods benchmarked using a masked sample $\hat{x}$ quantify the mutual information $I(\hat{x};c)$. Assuming the attribution mask $S$ operates as a patch that directly removes features from the input sample, this process inevitably allows the leakage of class-related information into the masked sample $\hat{x}$, leading to a leakage of $I(S;c)$. The leakage leads to the unfaithfulness of evaluation results by introducing class information from $S$ in $\hat{x}$, i.e., $I(\hat{x};c)\neq I(c;\hat{x}|S)$. To ensure benchmarking fidelity, the leaked features from $S$ should be eliminated, i.e., $I(S;c)\approx 0$. This relation is formalized by the below proposition.
@@
The proof of the proposition is provided in Appx. 10. In contrast to the prevalent existing benchmarks that assess attributions through feature removal, our benchmark recovers the features of the poisoned sample with those of the clean sample through the mask, i.e., $S\odot x\in x$. The recovered features which are part of the clean sample $x$ do not contain additional information related to the target class $\tilde{y}$. As a result, the leaked information from $S$ to class $\tilde{y}$ is negligible, if any. Thus, the evaluation process within our benchmark provides assurance of Attribution Verifiability. This also highlights an intrinsic superiority of our benchmark over the perturbation-based methods that are inadvertently influenced by feature leakage.
Overall, our benchmark offers the desired level of assurance for fulfilling the essential fidelity criteria, which is not achieved by the previous benchmarking techniques.
5 Consistent Attribution Evaluation
@@
Having established the foundations of a reliable benchmark for attribution methods, we further analyze the existing attribution techniques for their transparent benchmarking. Currently, a diverse range of attribution methods is available in the literature. It has been demonstrated that their evaluation results are significantly influenced by two common confounding factors, (1) post-processing techniques of attributions [26, 28] and (2) explained model outputs in attribution estimation [27]. However, to the best of our knowledge, no existing contribution explores a consistent evaluation paradigm for the different attribution methods. This section is aimed at establishing a fair paradigm for consistent benchmarking. To achieve that, we first discuss the attribution methods evaluated in this work, and their categories. Subsequently, our analysis is grounded in extensive experiments across diverse setups to establish a consistent setup [^1] Details of the models used in our experiments, as well as their comparisons, are reported in Appx. 11.2 and Appx. 12..
5.1 Attribution Methods and Categories
In our experiments, nine attribution methods are benchmarked including GCAM [32], FullGrad [10], Grad [6], GGCAM [32], SG [26], IG [8], IG-Uniform [33], AGI [34], and LPI [37]. We categorize these methods into three groups; namely CAM-based, gradient-based, and integration-based methods, based on their distinct underlying explanation processes, as explained below.
1) CAM-based Methods: Zhou et al. [31] proposed a class activation mapping (CAM) technique to localize the class-specific features of input samples. Assuming a feature map $A^{k}$ of the $k$-th convolutional layer before a softmax layer, CAM-based attribution methods $M^{c}_{CAM}$ (e.g., GCAM) calculate a weighted combination of activation maps $A^{k}$ for a class $c$ as
$M^{c}_{\text{CAM}}(x)=\Psi(ReLU(\sum_{k}w^{c}_{k}A^{k})),$ (8)
where $w^{c}_{k}=avg(\sum_{i}\sum_{j}\partial f_{c}(x)/\partial A^{k}_{i,j})$ indicates the activation weight by applying a global average pooling on the activation map $A^{k}$, and $\Psi$ indicates an interpolation operator to estimate an attribution map with the same scale as $x$ from a down-sampled activation map. In our experiments, we test two CAM-based attribution methods including GCAM and FullGrad. In contrast to CAM, GCAM avoid the use of additional structures enabling a generalized applications. FullGrad further integrates gradients of model biases to ensure the completeness axiom[8]. Since FullGrad retains a similar process of utilizing an interpolated activation map, we categorize it alongside GCAM as a CAM-based method. Although other CAM-based attribution methods also exist [49, 50], we test two representative ones in our experiments to cover this category.
2) Gradient-based Methods: Gradient-based attribution methods, such as Grad [6], employ a first-order Taylor expansion to approximate the non-linear model: $f_{c}(x)\approx w^{\intercal}*x+b$. Thus, the attribution of an input sample $x$ for a class $c$, can be computed using the gradients with respect to the input as
$M^{c}_{\text{Grad.}}(x)=\partial f_{c}(x)/\partial x.$ (9)
3) Integration-based Methods: Compared to gradient-based attribution methods, integration-based attribution methods integrate input gradients from a reference input $x^{\prime}$ to the feature $x$ along an integral path. Taking IG as an example, the attribution of an input feature $x_{i}$ for a class $c$ can be estimated as
$M^{c}_{i\ \text{Integr.}}(x,x^{\prime})=(x_{i}-x^{\prime}_{i})\times\int_{\alpha=0}^{1}\frac{\partial f_{c}(\tilde{x})}{\partial\tilde{x}_{i}}\bigg{|}_{\tilde{x}=x^{\prime}+\alpha(x-x^{\prime})}\mathrm{d}\alpha,$ (10)
where $\alpha$ indicates a linear integration path from the reference input $x^{\prime}$ to the input $x$, and $x^{\prime}$ is typically set as a zero vector in IG. Different integration-based attribution methods (IG-Uni, IG-SG [26], AGI, and LPI) are proposed to redefine the reference and the integral path. As these methods share the approach of estimating integrals to offer explanations, we categorize these methods as integration-based attribution methods. Due to their distinct theoretical guarantees in using different references [8, 37], we benchmark four integration-based attribution methods including IG, IG-Uni, AGI and LPI.
4) Other Methods: A variety of attribution methods also have been proposed for providing explanations through feature perturbation or removal (e.g., LIME [51], SHAP [52], occlusion [7], and mask [53]). However, these perturbation-based methods are revealed to exhibit unreliabilities [9, 8], as well as entailing a significant computational cost [38]. Therefore, we exclusively evaluate the propagation-based attribution methods. On the other hand, a few attribution methods, e.g., LRP [54] and DeepLift [9], fail to satisfy basic axioms like completeness and implementation invariance. Hence, these methods are not included in our evaluation.
@@ 5.2 Post-processing Choice
It is currently prevalent to use any of the absolute or original attribution scores (as computed by an attribution method) to explain model predictions, despite the fact that these two choices yield distinct explanations. In Figure 3, we provide visualizations of employing absolute and original attributions for different methods. There is a notable disparity in the highlighted regions of input samples between the two post-processing techniques. This underscores the susceptibility of visual outcomes to ambiguity, emphasizing the necessity of quantitative assessments to provide clarity and context. This problem is more pronounced for smaller input images due to the limited number of pixels. Although large ImageNet size images may yield more consistent attribution maps, it is crucial to recognize that their relative feature importances undergo shifts, often reflected in attributions alternating between negative and positive values. This begs for an enquiry into the right choice between the original and absolute values for benchmarking.
@@
In Fig. 4, we analyze the attack success rate and trigger recall for the two post-processing choices under BackX while fixing the image recovery equal to the trigger ratio. To facilitate comparison, we calculate the differences in benchmarking results between absolute and original values of attributions. The analysis is performed for CIFAR-10 [55], GTSRB [56] and ImageNet [57] datasets. For CAM-based methods, it can be observed that GCAM maintains consistent results across the datasets and our metrics [^2] While ReLU is typically applied to CAM-based results, we intentionally remove ReLU to preserve the original attributions.. FullGrad’s integration of gradients from biases leads to fluctuations among the datasets. In contrast, all three gradient-based methods consistently rely on taking absolute attribution values across all the datasets for improved performance. The results align with the finding of Smilkov et al. [26], suggesting that taking the absolute value of gradients can produce clearer attribution maps. One plausible explanation behind the insensitivity of gradient-based methods to the sign of attributions lies in their focus on capturing the magnitude of each feature’s influence on the model output using input gradients, irrespective of the direction of change in the feature’s value.
On the other hand, integration-based methods have a slightly more stable performance for different datasets. Their reliance on absolute values of attributions decreases as the mean value of the image pixels decreases. This can be attributed to the fact that these techniques accumulate gradients from a reference input. Images with lower mean pixel values tend to reduce the undesired gradient fluctuations due to their natural proximity to the reference. We draw the following observation from our analysis.
Observation 1 Basic CAM-based methods remain largely insensitive to post-processing techniques. Gradient-based attribution methods consistently depend on computing absolute attributions, whereas this reliance decreases in integration-based methods with a reduction in the mean pixel value.
Based on our observation, we choose the absolute attributions for gradient-based methods in the subsequent experiments. CAM-based and integration-based methods employ the original attributions to ensure the satisfaction of their well-known axioms [^3] FullGrad and integration-based methods are able to satisfy the completeness axiom with their original attributions..
Figure 5: The performance comparison of CAM-based, gradient-based and integrated-based attribution methods for output choice. (a) Difference between attack success rate when attributions are computed for softmax probabilities (prob.) and logits. (b) Trigger recall difference using softmax probabilities and logits.
[…]

### v2

@@ 3.2 Evaluation Metrics
In this part, we define metrics specifically tailored to our backdoor-based evaluation for BackX.
Attack Success Rate. Attack Success Rate (ASR) is a prevalent metric in backdoor attacks (Turner et al., 2019; Guo et al., 2022). It records the success of target prediction label $\tilde{y}$ for the input samples with true label $y$ when the trigger is embedded in the input samples. In our context, the ASR is formally defined as
@@
Here, the ASR is computed as an expectation over samples drawn from a dataset $\mathcal{D}$ through a boolean function $\bm{{\rm I}}[\cdot]$. In our benchmark, we leverage ASR to evaluate the ability of attribution methods to successfully identify the trigger with the recovered sample $\hat{x}$.
Trigger Recall. We then assess the localization capability of attribution methods. Trigger Recall (TR) is introduced to evaluate the ability of an attribution method to locate the embedded triggers. TR computes an expectation over poisoned samples $\tilde{x}$ by evaluating the alignment between its trigger mask $S^{*}_{\tilde{x}}$ and the attribution mask $S^{(k)}_{\tilde{x},\tilde{y}}$ of target class $\tilde{y}$ as
@@
where $S^{(k)}_{\tilde{x},\tilde{y}},S^{*}_{\tilde{x}}\in\{0,1\}^{n}$, and $k\in[0,1]$ is the fraction of the most important features identified.
[…]
@@
[…]
@@ 3.3 Benchmark Fidelity Examination
In this part, we closely examine how BackX guarantees the fidelity criteria defined in Sec. 2.1.
Functional Mapping Invariance & Metric Sensitivity. The BackX framework inherently ensures Functional Mapping Invariance by consistently employing a fixed Trojaned model $\tilde{f}$ throughout the benchmarking process, without introducing any perturbation $\zeta$ to the model parameters $\theta$. In addition, the proposed three metrics are designed to independently and precisely assess each component of the estimated attribution set ${\phi_{i}(x)}_{i=1}^{n}$, thereby guaranteeing Metric Sensitivity.
Input Distribution Invariance. To rigorously assess this criterion, we present Proposition 1 below.
@@
Let $f$ be a model that assigns the label $y$ to any input $x$ and to all inputs $\bar{x}$ within an $\epsilon$-ball under metric $\Omega$, i.e., $f(\bar{x})=f(x)$ whenever $\Omega(x,\bar{x})\leq\epsilon$. Suppose a Trojaned input is given by $\tilde{x}=x+v$, where $\Omega(x,\tilde{x})\leq\tilde{\epsilon}$ and $f(\tilde{x})\neq f(x)$ due to the trigger $v$. Assume $\hat{x}$ is a reconstruction of $x$ from $\tilde{x}$ such that $\Omega(x,\hat{x})\leq\tilde{\epsilon}$. If $\tilde{\epsilon}\leq\epsilon$, then $\hat{x}$ lies within the $\epsilon$-ball of $x$, and thus $f(\hat{x})=f(x)$.
The proof of Proposition 1 is in App. A.3. It describes when the input undergoes slight changes due to neural Trojans, the restored sample $\hat{x}$ remains within a permissible boundary, ensuring it stays within the original distribution, provided the shift $\tilde{\epsilon}$ is sufficiently small compared to $\epsilon$.
@@
Given the stealthy nature of neuron backdoors, the constraint $\tilde{\epsilon}$ is negligibly small for well-trained models (Szegedy et al., 2013), which implies $\Omega(x,\hat{x})\leq\Omega(x,\bar{x})$. This suggests that $\tilde{\epsilon}\leq\epsilon$, ensuring that the reconstructed sample $\hat{x}$ remains largely within the original data distribution. Whereas we acknowledge that BackX cannot perfectly eliminate the influence of input distribution shift, the Input Distribution Invariance is only subtly disturbed and remains effectively bounded by $\epsilon$ throughout benchmarking.
Attribution Verifiability. BackX focuses on benchmarking using recovered samples, which may unintentionally reveal information about the attribution mask used in recovery (Rong et al., 2022), thereby compromising attribution ground truth fidelity. To ensure Attribution Verifiability, it is crucial to rigorously assess information leakage from the recovery samples.
@@
We commence by examining the entropy of a singular variable $x$, quantifying the amount of information as $H(x)=-\sum_{x_{i}\in x}P(x_{i}){\rm log}P(x_{i})$. The classification performance is correlated with the mutual information $I(x;c)$ between the input $x$ and a class $c$. Attribution methods benchmarked using a masked sample $\hat{x}$ quantify the mutual information $I(\hat{x};c)$. Assuming the attribution mask $S$ operates as a patch that directly removes features from the input sample, this process inevitably allows the leakage of class-related information into the masked sample $\hat{x}$, leading to a leakage of $I(S;c)$. The leakage leads to unfaithful evaluations by introducing class information from $S$ in $\hat{x}$, i.e., $I(\hat{x};c)\neq I(c;\hat{x}|S)$. To ensure benchmarking fidelity, the leaked features from $S$ should be eliminated, i.e., $I(S;c)\approx 0$. The below proposition formalizes this relation.
@@
The proof of the proposition is provided in App. A.3. In contrast to the existing benchmarks using feature removal, our benchmark recovers the features of the poisoned sample with those of the clean sample through the mask. The recovered features which are part of the clean sample $x$ do not contain additional information related to the target class $\tilde{y}$. As a result, the leaked information from $S$ to class $\tilde{y}$ is negligible, if any. Thus, the evaluation process within our benchmark provides assurance of Attribution Verifiability. In App. A.5, further fidelity examination of Attribution Verifiability for the used models and metrics is provided.
Thus, we theoretically show that it is not the mere existence of neural Trojans that yields benchmarking fidelity, but that BackX’s tailored strategies with tighter theoretical guarantees establish the backdoor-based route as a definitive pathway to fidelity. A further discussion is provided in App. A.13. Moreover, we formally establish that BackX offers the desired level of assurance for fulfilling the essential fidelity criteria, which is not achieved by the prior benchmarking techniques.
4 Standardized Attribution Evaluation
@@
Having established the foundations of a reliable benchmark for attribution methods, we further analyze the existing attribution techniques for their transparent benchmarking. Currently, a diverse range of attribution methods is available in the literature. It has been demonstrated that their evaluation results are significantly influenced by two common confounding factors, (1) post-processing techniques of attributions (Smilkov et al., 2017; Yang et al., 2023b) and (2) explained model outputs in attribution estimation (Wang & Wang, 2022). However, to the best of our knowledge, no existing contribution explores a consistent evaluation paradigm for the different attribution methods. Our empirical analysis below is aimed at establishing a fair paradigm for consistent benchmarking.
To achieve our goal, we analyze diverse attribution methods including Grad-CAM (GCAM) (Selvaraju et al., 2017), FullGrad (Srinivas & Fleuret, 2019), Input Gradients (Grad) (Simonyan et al., 2014), Guided GCAM (GGCAM) (Selvaraju et al., 2017), SmoothGrad (SG) (Smilkov et al., 2017), Integrated Gradients (IG) (Sundararajan et al., 2017), IG-Uniform (Sturmfels et al., 2020), AGI (Pan et al., 2021), and LPI (Yang et al., 2023a). We categorize these techniques into three groups based on their distinct methodologies, namely; CAM-based methods (GCAM and FullGrad), gradient-based methods (Grad, GGCAM and SG), and integration-based methods (IG, IG-Uni, AGI and LPI). Further discussion and results on attribution methods such as LRP, LIME, and others (Ribeiro et al., 2016; Shrikumar et al., 2017; Novello et al., 2022) are presented in App.A.6 and App.A.10.2.
@@ 4.1 Post-processing Choice
It is currently prevalent to use any of the absolute or original attribution scores to explain model predictions, despite the fact that these two choices yield distinct explanations. This begs for an enquiry into the right choice between the original and absolute values for benchmarking. In Fig. 3, we analyze the attack success rate and trigger recall for the two post-processing choices under BackX while fixing the image recovery equal to the trigger ratio. To facilitate comparison, we calculate the differences in benchmarking results between absolute and original values of attributions. The analysis is performed for CIFAR-10 (Krizhevsky et al., 2009), GTSRB (Houben et al., 2013) and ImageNet (Russakovsky et al., 2015) datasets. For CAM-based methods, GCAM maintains consistent results across the datasets and our metrics. FullGrad’s integration of gradients from biases leads to fluctuations among the datasets. In contrast, all three gradient-based methods consistently rely on taking absolute attribution values across all the datasets for improved performance. One plausible explanation behind the insensitivity of gradient-based methods to the sign of attributions lies in their focus on capturing the magnitude of each feature’s influence on the model output using input gradients, irrespective of the direction of change in the feature’s value.
On the other hand, integration-based methods have a slightly more stable performance across data sets. Their reliance on absolute attribution scores decreases as the mean value of the image pixels decreases. This can be attributed to the fact that these techniques accumulate gradients from a reference input. Images with lower mean pixel values tend to reduce the undesired gradient fluctuations due to their natural proximity to the reference. We draw the following observation from our analysis.
Observation 1. Basic CAM-based methods remain insensitive to post-processing techniques. Gradient-based methods consistently depend on computing absolute attributions, whereas this reliance decreases in integration-based methods with a reduction in the mean pixel value.
Based on our observation, we use absolute attributions for gradient-based methods in the subsequent experiments. CAM- and integration-based methods use original attributions to ensure their axioms.
@@
Figure 4: The performance comparison of attribution methods for output choice. (a) Difference between attack success rate when attributions are computed for softmax probabilities (prob.) and logits. (b) Trigger recall difference using softmax probabilities and logits.